OpenStudy (anonymous):

PLZ HELP!! darla is building a new desk. to make sure she had made a square corner she measures 9ft. from the corner along one edge and 6ft. from the corner along the other edge. how long should the diagonal be between those two points if the corner is a right angle? (round to the nearest thousandth)

OpenStudy (anonymous):

Pythagorean Theorem for this one: a^2+b^2=c^2 \[6^{2}+9^{2}=c^{2}\] \[117=c^{2}\] \[\sqrt{117}=c\] The sqrt of 117 is 10.817